Hi again,
I could use some assistance with this problem:
Prove that if the sum of two numbers is constant, then their product is maximum if the numbers are equal.
Thank you,
Louis

Answer by stanbon(75887)

Prove that if the sum of two numbers is constant, then their product is maximum if the numbers are equal.
1st: Let x+y = k, where k is some constant.
2nd: Then y=k-x
3rd: The product of x and y is xy=x(k-x)=kx-x^2
This is a quadratic with a=-1, b=k
The maximum value of this product occurs when x=-b/2a
x=-b/2a= -k/-2 = k/2
Substitute that into the 1st equation to get the following:
(k/2)+ y=k
y=k/2
So both x and y are k/2; therefore they are equal.
Cheers,
Stan H.